OpenStudy (anonymous):

write an equation in point-slope form of the line that passes through the given points (7,2), (2,12)?

OpenStudy (anonymous):

first find the slope m = (12-2)/(2-7) = 10/-5 = -2 then use point-slope y - y1 = m(x - x1) or y -12 = -2(x - 2) distribute y - 12 = -2x + 4 add 12 y = -2x + 16